# R2R: Efficiently Navigating Divergent Reasoning Paths with Small-Large Model Token Routing

### v1

@@ 5.4 Ablation Study
Starting from R2R in the first row of Table 4, we evaluate the effectiveness of our design by retraining the router with alternative objectives or reduced inputs, keeping all other settings fixed. All experiments are conducted on the AIME benchmark.
Routing objective. As discussed in Section 3, we categorize different next-token predictions as either neutral or divergent. R2R improves efficiency by tolerating neutral differences and only routing truly divergent tokens to the LLM. When the router is trained to use the LLM for all different tokens, it fails to reach the original accuracy within the same amount of LLM usage, facing 1.4×1.4$\times$1.4 × accuracy degradation, as shown in the second row of Table 4. This confirms that restricting LLM usage to only divergent tokens is crucial for reducing cost while maintaining high accuracy.
Router Input. As discussed in Section 4, both SLM logits and token embeddings are strong indicators of divergence and are used as router inputs. When these features are gradually removed, routing accuracy drops by up to 1.3×1.3$\times$1.3 ×, underscoring their importance. Note that while SLM logits can, in principle, be computed from last-layer hidden states within the router. However, doing so requires the capacity of the 234M-parameter embedding layer, which exceeds the capacity of the 56M-parameter neural router.
Table 4: Ablation study on routing objectives and router inputs. HS denotes last-layer hidden states; Token denotes token embedding. Italicized words indicate ablation focuses.
Objective Router Input Acc. #Token(K) LLM Rate(%) Param.(B) Cost(K×$\times$×B) Latency(s)
Divergent HS+Token+Logits 55% 18.4 12.4 5.5 101 218
Different HS+Token+Logits 40% 19.7 14.4 6.1 120 333
Divergent HS+Token 47% 18.8 13.3 5.8 109 353
HS 42% 18.4 14.1 6.0 110 345
@@ 5.5 Routing Result Observation
Following prior work [21], we further segment the thinking process into sequential thoughts based on tokens such as Wait and Alternatively. Figure 6(b) examines the LLM usage ratio within each thought. It shows that R2R relies more on the LLM at the beginning and end of each thought. This aligns with the expectation that the initial tokens set the direction for the thought, while the concluding tokens determine whether to end the thought, branch into alternatives, or continue deeper reasoning. Notably, these routing patterns are not hand-crafted but naturally emerge from router training. It helps R2R to effectively allocate LLMs for more better test-time scaling.
@@ 6 Conclusion
In this paper, we introduce R2R, a token-level routing method that enables SLMs to follow LLM reasoning paths by correcting only divergent tokens. We design a path-following routing strategy for effective data labeling and investigate predictive indicators for a neural router to achieve accurate token selection. Experiments on challenging benchmarks demonstrate that R2R outperforms the R1-14B model with less than 7B average parameters, boosts SLM performance by 4.6$\times$ with under 15% LLM usage, and achieves a 2.8$\times$ wall-clock speedup over the LLM at comparable accuracy. Limitations. Our current routing strategy assumes greedy sampling. Exploring alternative sampling strategies may further enhance versatility. Additionally, more dedicated system-level optimizations are required to fully realize the theoretical cost benefits of R2R.
@@
This work was supported by National Natural Science Foundation of China (No. 62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164, 92364201), Tsinghua EE Xilinx AI Research Fund, and Beijing National Research Center for Information Science and Technology (BNRist). We thank Zinan Lin, Xuefei Ning, and Donglin Yang for their valuable discussions and suggestions. We thank Chao Xiong for his support with the SGLang interface. We also thank all the support from Infinigence-AI.
@@ A.1 Router achitecture
Neural network backbone. For the router architecture, we adopt a six-layer feed-forward network (FFN) with residual connections between blocks as the backbone, using a hidden size of 1024. The architecture of each FFN follows the common design used in previous LLMs [38]. Each block begins with LayerNorm for input normalization, followed by a pair of linear projections forming an expand-then-contract structure with an expansion factor of 4. Dropout is applied to each linear layer, with a GELU activation function between them. These blocks are connected using residual connections. At the end of the last block, we apply an additional layer normalization and a linear layer to convert the output to a single value, followed by a sigmoid function to produce the normalized prediction of the router. A predefined threshold $p_{\text{th}}$ between 0 and 1 is used for generating binary results from the router output. Predictions above the $p_{\text{th}}$ are considered that current tokens diverge from the LLM’s reasoning path.
@@ A.2 Routing Data
Training dataset. Our training data for the router are sourced from tasks across three distinct scenarios: mathematics, code, and question answering (QA). The mathematics problems are drawn from the American Invitational Mathematics Examination (AIME) [10], covering the years 1983 to 2022. Code and QA problems are sampled from Bespoke-Stratos-17k dataset [31]. We use only the formatted questions from these datasets as prompts and generate responses using DeepSeek-R1-Distill-Qwen-32B, with the temperature set to 0 and a maximum generation length of 32,768 tokens. Only responses that contain an end-of-sequence (EOS) token within the specified length are retained as effective samples, which will be used for subsequent stages of our data generation pipeline, as discussed in Section 4.1.
Validation dataset. Our validation dataset are constructed in the exact same way as the training data, but with different queries. The validation dataset comprises all 30 problems from AIME 2023, 69 coding problems from the Bespoke-Stratos-17k dataset that are excluded from the training set, and 60 QA problems selected from the GPQA-Extended [35] dataset.
@@ Appendix B Additional Experiment Results
B.1 Speculative Decoding
Table 5: Comparison of latency and average speed across methods and batch sizes. Subscripts note the standard deviations across AIME. To ensure measurement accuracy, each query was replicated according to its respective batch size. The Speed(tok/s) denotes tokens processed per second for a single batch operation.
Speed(tok/s) Latency(s)
Method batch = 1 batch = 2 batch = 4 batch = 1 batch = 2 batch = 4
Eagle2 71.4 56.5 38.3 244 $\pm$194 300 $\pm$257 459 $\pm$388
HASS 73.3 60.4 37.7 256$\pm$197 255 $\pm$246 474 $\pm$417
Experiment hyperparameters. For both Eagle2 and HASS, we follow their official implementations to set the total number of draft tokens to 60, with a draft tree depth of 6, and select 10 nodes during the expansion phase.
Limitations of speculative decoding. Speculative decoding accelerates inference by fixing the low parallelism of LLM decoding. It drafts outputs through SLM sequential decoding, then periodically verifies them with high-parallel LLM prefilling. However, in the official implementation, the prefill length (60) of LLM verification substantially exceeds the depth of the draft tree (6), significantly increasing computational demands. At batch size 1, the overall process remains memory bandwidth bound, meaning that the excessive computation required for prefilling does not significantly impact inference latency. As batch size increases, the process transitions to being computation bound, resulting in noticeable latency increases as shown in Table 5.
B.2 Data Generation Overhead
@@
As Table 6 shows, our four-stage pipeline completes in 56 h (448 GPU hours). The LLM response stage dominates runtime (35 h, 280 GPU hours), but it can be mitigated by directly utilizing the responses of LLMs from open-source SFT datasets, provided they were generated by the same LLM used for routing. The SLM prefill step is highly efficient, requiring only 0.1 h of wall clock time. The subsequent LLM continuation and verification stages take 7 h (56 GPU hours) and 14 h (112 GPU hours), respectively. Compared to downstream tasks, the overall data generation pipeline remains relatively lightweight and efficient.
B.3 Performance-Cost Trade-off
B.4 Influence of Routing Threshold

### v2

@@ 5.4 Ablation Study
Starting from R2R in the first row of Table 4, we evaluate the effectiveness of our design by retraining the router with alternative objectives or reduced inputs. We adjust the routing thresholds $p_{\text{th}}$ of ablation variants to match or slightly exceed the LLM rate of our original router. All experiments are conducted on the AIME benchmark with all other settings held constant. Further discussions on the trade-off between the LLM usage rate and recall are provided in Appendix B.5.3.
Routing objective. As discussed in Section 3, we categorize different next-token predictions as either neutral or divergent. R2R improves efficiency by tolerating neutral differences and only routing truly divergent tokens to the LLM. When the router is trained to use the LLM for all different tokens, it fails to reach the original accuracy within the same amount of LLM usage, facing $1.4\times$ accuracy degradation, as shown in the second row of Table 4. This confirms that restricting LLM usage to only divergent tokens is crucial for reducing cost while maintaining high accuracy.
Router input. As discussed in Section 4, both SLM logits and token embeddings are strong indicators of divergence to be used as router inputs. When gradually remove these features, accuracy drops by up to $1.3\times$, underscoring their importance. Note that while SLM logits can be computed from last-layer hidden states within the router in principle. However, doing so requires the capacity of the 234M-parameter embedding layer, which exceeds the capacity of the 56M-parameter neural router.
Table 4: Ablation study on routing objectives and router inputs. HS denotes last-layer hidden states; Token denotes token embedding. Recall denotes the recall rate of divergent token on the validation dataset. Italicized words indicate ablation focuses.
Objective Router Input Acc. Recall LLM Rate #Token(K) Param.(B) Cost(KB) Latency(s)
Divergent HS+Token+Logits 55% 95% 12.4% 18.4 5.5 101 218
Different HS+Token+Logits 40% 88% 13.1% 21.0 5.7 119 228
Divergent HS+Token 47% 85% 13.3% 18.8 5.8 109 253
HS 42% 83% 14.1% 18.4 6.0 110 245
SLM–LLM combination. We extend R2R to the Qwen3 family by fixing the LLM and varying the SLM to form three combinations: Qwen3-0.6B+8B, Qwen3-1.7B+8B, and Qwen3-4B+8B. As Figure 10 in Appendix B.5.3 shows, with the LLM fixed, smaller SLM delivers a better Pareto frontier. It indicates that, when selecting SLM–LLM combinations, favoring a smaller SLM is often more efficient for the same accuracy target due to lower cost. We further analyze R2R alongside orthogonal techniques such as MoE and query-level routing in Appendix B.4.
@@ 5.5 Routing Result Observation
Following prior work [wang2025overthink], we further segment the thinking process into sequential thoughts based on tokens such as Wait and Alternatively. Figure 6(b) examines the LLM usage ratio within each thought. It shows that R2R relies more on the LLM at the beginning and end of each thought. This aligns with the expectation that the initial tokens set the direction for the thought, while the concluding tokens determine whether to end the thought, branch into alternatives, or continue deeper reasoning. Notably, these routing patterns are not hand-crafted but naturally emerge from router training. It helps R2R to effectively allocate LLMs for more better test-time scaling.
@@ 6 Conclusion
[…]
Limitations. Our routing strategy is chiefly tuned and tested for greedy sampling. While we have validated R2R on a limited set of alternative sampling policies, broader exploration could improve versatility. Further system-level optimizations are also needed to realize its full cost benefits.
@@
This research was supported by the National Natural Science Foundation of China (No. 62506197, 62203257, 62325405, 62031017, 62406159), Tsinghua University Initiative Scientific Research Program, Tsinghua-Efort Joint Research Center for EAI Computation and Perception, Beijing National Research Center for Information Science, Technology (BNRist), Beijing Innovation Center for Future Chips, and State Key laboratory of Space Network and Communications.
@@ A.1 Router Architecture
Neural network backbone. For the router architecture, we adopt a six-layer feed-forward network (FFN) with residual connections between blocks as the backbone, using a hidden size of 1024. The architecture of each FFN follows the common design used in previous LLMs [zhang2022opt]. Each block begins with LayerNorm for input normalization, followed by a pair of linear projections forming an expand-then-contract structure with an expansion factor of 4. Dropout is applied to each linear layer, with a GELU activation function between them. These blocks are connected using residual connections. At the end of the last block, we apply an additional layer normalization and a linear layer to convert the output to a single value, followed by a sigmoid function to produce the normalized prediction of the router. A predefined threshold $p_{\text{th}}$ between 0 and 1 is used for generating binary results from the router output. Predictions above the $p_{\text{th}}$ are considered that current tokens diverge from the LLM’s reasoning path.
@@ A.2 Routing Data
Training dataset. Our training data for the router are sourced from tasks across three distinct scenarios: mathematics, code, and question answering (QA). The mathematics problems are drawn from the American Invitational Mathematics Examination (AIME) [aimedata], covering the years 1983 to 2022. Code and QA problems are sampled from Bespoke-Stratos-17k dataset [bespokestratos]. We use only the formatted questions from these datasets as prompts and generate responses using DeepSeek-R1-Distill-Qwen-32B, with the temperature set to 0 and a maximum generation length of 32,768 tokens. Only responses that contain an end-of-sequence (EOS) token within the specified length are retained as effective samples, which will be used for subsequent stages of our data generation pipeline, as discussed in Section 4.1.
Validation dataset. Our validation dataset are constructed in the exact same way as the training data, but with different queries. The validation dataset comprises all 30 problems from AIME 2023, 69 coding problems from the Bespoke-Stratos-17k dataset that are excluded from the training set, and 60 QA problems selected from the GPQA-Extended [rein2023gpqa] dataset.
@@ Appendix B Additional Experiment Results
B.1 Verifier Robustness
B.1.1 LLM Verifier Against Human Verifier
Setup. We conducted a human evaluation to validate the verifier’s reliability. Specifically, we recruited four undergraduate students to independently label 1,357 differences as either neutral or divergent. The differences are between R1-1.5B and R1-32B on the first six AIME-2024 questions. Three annotators’ labels determined the ground truth for divergence by majority voting (general divergence, positive rate: 24.8%) and unanimous consent (core divergence, positive rate: 11.9%). The fourth annotator’s labels, along with several LLM verifiers, were then compared against these ground truths. All annotators and LLM verifiers received identical contexts and instructions.
Results and discussion. Our selected verifier (Qwen2.5-72B) closely matches human expert performance. The detailed results are shown in Table 5.
Table 5: Verifier performance against majority vote (General Divergence, Positive Rate: 24.8%) and unanimous consent (Core Divergence, Positive Rate: 11.9%) ground truths.
Verifier Majority Voting (General) Unanimous Consent (Core) Positive Rate
Accuracy Recall Precision Accuracy Recall Precision
Human 0.82 0.84 0.59 0.76 0.96 0.32 0.35
Qwen2.5-72B 0.84 0.88 0.62 0.76 0.97 0.33 0.35
Qwen2.5-7B 0.78 0.85 0.54 0.71 0.94 0.28 0.39
Qwen2.5-3B 0.75 0.69 0.50 0.72 0.75 0.26 0.34
B.1.2 Impact of Verifier Quality
Setup. Next, we examined the sensitivity of our method to verifier quality. We applied the sentence-level path-following routing method (Equations 35) to the AIME 2024 benchmark across verifiers (Qwen2.5-3B to Qwen2.5-72B) within 8K token budget.
Results and Discussion. Our analysis reveals that even moderate-recall verifiers (e.g., Qwen2.5-3B) still significantly improve reasoning path guidance compared to the distilled 7B baseline. Note that the current verifier prompt is tuned for the Qwen2.5-72B verifier. Empirically, smaller verifiers should benefit from prompts biased towards classifying ambiguous cases as divergent, thereby enhancing recall.
Table 6: Impact of verifier quality on path-following routing performance on AIME 2024.
Type Model Param. (B) Labeling Verifier AIME’24 #Acc.@8K
LLM Baseline 32 - 17
SLM Baseline 1.5 - 2
Distilled 7 - 8
Path-following 2.3 Qwen2.5-72B 16
Path-following 2.5 Qwen2.5-7B 12
Path-following 2.2 Qwen2.5-3B 11
B.2 Generalizability
B.2.1 Data Labeling Method Generalizability
The labeling method of R2R, relying on semantic comparison rather than formal verification, can easily generalize across tasks and achieve high performance even in some out-domain tasks. Unlike methods requiring task-specific external tools (e.g., formal proofs  [xin2024deepseekproverv15harnessingproofassistant], code execution  [jung2025codeexecutiongroundedsupervision]), R2R uses an LLM verifier to identify general semantic divergence in meaning, reasoning, logic, or conclusions F.1.
In practice, we apply the identical labeling strategy and verifier prompt across closed-form math, coding tasks, and open-ended QA tasks. This consistency enables our router to naturally generalize to unseen tasks during inference, as demonstrated later.
For tasks with subjective divergence criteria that are challenging to identify semantically within a single sentence, the continuation length can be extended. This adjustment gradually transitions from sentence-level routing to full routing, the latter of which only requires overall response quality evaluation, which is a feasible requirement commonly met for LLM tasks.
B.2.2 Router Generalizability
Generalize across benchmarks. Our router leverages outputs from the SLM (e.g., logits) to predict token divergence. Benefiting from the inherent generalizability of SLMs, indicators such as logits entropy robustly identify divergent tokens across different tasks. To validate the router’s generalizability, we directly apply our router, trained for math, QA, and code, to additional benchmarks: Arena-Hard for Dialog  [li2024crowdsourceddatahighqualitybenchmarks], and the Philosophy split from MMLU-Redux [gema2025mmlu], which are never included in our training dataset. As table 7 shows, although the cost–accuracy trade-off gains are a bit less pronounced on out-of-domain tasks, R2R nonetheless exhibits strong generalization, which continues to outperform the 14B model with less than 7B parameter size.
Table 7: Performance of R2R on Arena-Hard for Dialog and MMLU-Redux-Philosophy.
Dialog Philosophy
Type Model(s) Score Param. Cost Acc. Param. Cost
LLM R1-32B 5.0 32 65.9 79 % 32 32.2
SLM R1-1.5B 0.2 1.5 14.3 38 % 1.5 8.7
Distill R1-7B 0.3 7 35.0 57 % 7 20.6
Distill R1-14B 2.2 14 56.5 77 % 14 18.5
R2R Ours 2.8 6.1 40.9 81 % 6.7 8.3
Generalize to different LLMs. We test whether a router, trained on data from a specific SLM-LLM pair (e.g., 0.6B-32B), can be effectively applied to a different pair (e.g., 0.6B-8B) without retraining. The experiments are conducted with the Qwen3 model family (0.6B, 8B, and 32B), tested on the AIME (2024+2025) benchmark with a target recall of 0.95 for divergent tokens. The results are summarized in Table 8. Our empirical results indicate acceptable performance when directly substituting the LLM of a trained router without additional retraining. This can be attributed to the common divergence patterns shared by strong LLMs paired with the same weaker SLM.
Table 8: Router generalizability on the AIME benchmark. Param. denotes the average activated parameters per token in billions; Cost is the average output tokens (thousands) per query multiplied by average parameters (billions) .
Type Model(s) Router Variant Acc. Param. Cost
SLM Qwen3-0.6B - 12% 0.6 15
LLM Qwen3-8B - 67% 8.0 130
LLM Qwen3-32B - 75% 32.0 469
R2R 0.6B+8B Trained on 0.6B+8B 65% 2.7 49
Generalized from 0.6B+32B 53% 2.6 50
R2R 0.6B+32B Trained on 0.6B+32B 67% 10.2 154
Generalized from 0.6B+8B 70% 9.9 151
B.3 Efficiency Analysis
B.3.1 Data Generation Overhead
@@
As Table 9 shows, our four‑stage pipeline completes in 56 h (448 GPU hours). The LLM response stage dominates runtime (35 h, 280 GPU hours), but it can be mitigated by directly utilizing the responses of LLMs from open‑source SFT datasets, provided they were generated by the same LLM used for routing. The SLM prefill step is highly efficient, requiring only 0.1 h of wall clock time. The subsequent LLM continuation and verification stages take 7 h (56 GPU hours) and 14 h (112 GPU hours), respectively. Compared to downstream tasks, the overall data generation pipeline remains relatively lightweight and efficient.
B.3.2 Computation and Memory Access
Following prior work [hoffmann2022training, yang2024glitches, wenheng2025cdllm], we analyze the computational and memory access overhead of R2R against various baselines on the AIME benchmark. Table 10 presents a detailed comparison of total computation (TFLOPs), total memory access (TB), and the average memory access per token (GB).
Table 10: Computational and memory overhead on the AIME benchmark. R2R demonstrates a significant reduction in memory access compared to the large model (R1-32B) and speculative decoding methods, with only a marginal increase in computation.
Model
Total Computation
Total Memory
Access (TB)
Avg. Memory Access
Per Token (GB)
R1-1.5B 157 104 3.7
R1-14B 630 503 29.4
R1-32B 1136 966 63.6
Eagle2 25490 515 29.6
HASS 26809 544 28.9
R2R 1502 216 11.7
Given that the decoding process in large language models is predominantly memory-bound, the volume of memory access is a critical factor influencing overall throughput. As shown in the table, our method, R2R, markedly reduces total memory access to just 216 TB, a 4.5x reduction compared to the R1-32B model. This efficiency is achieved with only a modest increase in total computation, which primarily arises from the additional KV-cache updates required when switching between the LLM and SLM.
In contrast, speculative decoding approaches such as Eagle2 and HASS incur substantially higher computational costs—more than 22x that of R1-32B. This overhead is a direct consequence of their complex tree-structured draft and verification processes, which can involve up to 60 tokens per cycle. Despite this significant computational demand, these methods also achieve considerable throughput improvements by effectively reducing the memory access bottleneck during the decoding phase. R2R, however, provides a more balanced trade-off, achieving significant memory savings without the extreme computational overhead of speculative methods.
B.3.3 Inference Latency Breakdown
We integrated R2R into the production-level SGLang serving infrastructure, which is demonstrated by the high absolute throughput results for our method and the baselines. To assess the latency implications of our token-level routing, we performed a detailed runtime breakdown. The analysis was conducted on the AIME benchmark, averaged over 60 questions, using two A800 GPUs.
The results, shown in Table 11, indicate that the token-level routing decisions made by our 56M parameter router introduce minimal latency overhead (5.96%). Given the small size of the neural router, we anticipate that further runtime improvements can be achieved through additional system-level optimizations.
Table 11: Runtime breakdown of the R2R system components on the AIME benchmark.
Component Percentage of Total Time
Router 5.96%
LLM (R1-32B) 64.97%
SLM (R1-1.5B) 26.77%
Others 2.30%
B.3.4 LLM Activation Intervals
The cost efficiency of the routing system largely depends on how frequently the LLM is activated during generation. Since the EXTEND operation of the LLM is inherently compute-bounded when consecutive LLM calls are spaced further apart. In other words, given a fixed LLM activation rate, longer intervals between two LLM invocations lead to better GPU utilization and less scheduling overhead per generated token.
To better understand this effect, we analyze the distribution of the interval between consecutive divergent tokens. Figure 7 illustrates these distributions for both the training dataset and actual inference traces on AIME25. As shown, the actual distribution closely aligns with that observed in training dataset, suggesting that the router generalizes well to real inference scenarios. More importantly, both distributions exhibit a long-tailed shape, indicating that divergence events occur sparsely and irregularly, which allows R2R to sustain high computational efficiency without overusing the LLM.
Figure 7: (a) Divergent token interval distribution for training dataset; (b) Reference token interval distribution on AIME25
B.4 Comparison and Compatibility with Orthogonal Techniques
B.4.1 R2R and Mixture-of-Experts (MoE)
We compare our method, R2R, with the Mixture-of-Experts (MoE) architecture, another prominent approach for efficient model scaling. This section begins with a conceptual comparison, followed by empirical results, and concludes by discussing how the two methods can be integrated to create a more favorable performance-efficiency trade-off.
Conceptual comparison. As noted, both MoE and R2R leverage partial activation, but they differ in key design aspects, as summarized in Table 12. This conceptual distinction motivates their potential integration, which we explore next.
Table 12: Conceptual comparison between MoE and R2R.
Design Aspect MoE R2R
Partial activation Yes Yes
Routing granularity Fine-grained (Experts) Coarse-grained (Models)
Subject model sizes Equal parameters per expert Different parameters per model
Training overhead Full training from scratch Router training only
Supervision Next-token predictions Explicit routing decisions
Empirical comparison. We evaluate the MoE model Qwen3-30BA3B and an R2R model (routing between Qwen3-0.6B and Qwen3-8B) on the AIME benchmark. The 60 questions are split into Easy (28) and Hard (32) subsets based on whether the question ID is less than or equal to 7. The results in Table 13 highlight the complementary strengths of each approach. Thanks to extensive pretraining, MoE models can match dense model performance on challenging tasks. However, their fixed per-token overhead limits efficiency on simpler inputs. In contrast, R2R dynamically adjusts computation based on token-level divergence, achieving lower cost on easier examples, though its lighter training cannot match MoE performance on harder questions given the same activated parameters. Given their complementary advantages, we explore the integration of both methods.
Table 13: Empirical comparison of R2R and MoE on the AIME benchmark, split by difficulty.
Type Model(s) Easy Hard
Acc. Param. Acc. Param.
R2R Qwen3-0.6B + Qwen3-8B 93% 2.9 41% 2.6
MoE Qwen3-30BA3B 93% 3.3 59% 3.3
MoE for R2R. R2R is directly compatible with efficiency-optimized models like MoEs. We validate R2R’s performance using the Qwen3-30BA3B MoE as the LLM, as shown in Table 14. Combining R2R with an MoE model significantly improves the Pareto frontier, achieving robust accuracy at extremely low average activated parameters.
Table 14: Performance of R2R when using an MoE model as the LLM.
Type Model Acc. Param. Cost
Dense Qwen3-0.6B 12% 0.6 15
Dense Qwen3-1.7B 37% 1.7 33
Dense Qwen3-32B 75% 32.0 469
R2R 0.6B + 32B 67% 10.2 154
MoE Qwen3-30BA3B 75% 3.0 51
R2R 0.6B + 30BA3B 68% 1.3 22
R2R for MoE. Conversely, R2R’s principles can enhance future MoE designs. Current MoE methods generally use uniform-sized experts; introducing mixed-sized experts could enable query-adaptive activation. Furthermore, supplementing an MoE’s training objective with explicit divergence or difficulty-based routing supervision from R2R may encourage more adaptive usage of larger experts for only the most critical decoding steps. Although retraining MoE models is beyond this paper’s scope, this presents an exciting direction for future research.
B.4.2 R2R and Query-level Routing (QR)
Query-level routing (QR) is methodologically orthogonal to R2R. Consequently, the composition of the two is a natural and well-motivated design choice. We investigated combining R2R with query-level routing methods using the Qwen3 model series with varied sizes, evaluated on the AIME benchmark. Specifically, we treated R2R-S (0.6B+8B) and R2R-L (0.6B+32B) as the SLM and LLM, respectively, for query-level routing.
Table 15: Performance of R2R when combining R2R and query-level routing (QR).
Type Model Acc. Param. (B) Cost (KB)
SLM Qwen3-0.6B 12% 0.6 15
LLM Qwen3-8B 67% 8 130
LLM Qwen3-32B 75% 32 469
R2R-S Qwen3-0.6B + 8B 59% 2.1 40
R2R-L Qwen3-0.6B + 32B 67% 10.2 154
R2R-M Qwen3-4B + 8B 65% 5.4 95
QR-BERT Qwen3-0.6B + 32B 25% 5.4 230
QR-LLM Qwen3-0.6B + 32B 23% 5.4 238
QR-MF Qwen3-0.6B + 32B 32% 5.4 216
QR-SW Qwen3-0.6B + 32B 30% 5.4 220
QR-BERT R2R-S + R2R-L 63% 5.4 179
QR-LLM R2R-S + R2R-L 67% 5.4 175
QR-MF R2R-S + R2R-L 58% 5.4 181
QR-SW R2R-S + R2R-L 62% 5.4 183
From Table 15, we can conclude that (1) R2R models generally serve as superior SLM/LLM candidates for query-level routing compared to original single-model setups, consistently advancing the Pareto frontier; (2) The interaction between medium-sized R2R models and query-level routing involving smaller or larger R2R-LLMs is empirically non-trivial. This observation suggests exciting opportunities for future research into sophisticated routing strategies, which combine the search space of query and token-level routing methods.
B.5 Additional Discussions and Analysis
B.5.1 Influence of Routing Threshold